ds317
General notes
- A few status is slewing errors for all three telescopes. (Arwin)
- Eremote controls not connecting for ops8, running fine on ops4. (Arwin)
Hobart 12m
Disk VSN: HOB+0120
Volume at beginning: 0.000 GB
- 01:23UT - Tracking started after disk change and scheduler issues (Ross)
- 09:20UT Telescope removed from dynamic schedule by Jamie and Lucia (Bryn)
- 10:30UT Telescope added back to dynamic schedule (Bryn)
- 10:34UT +1day Telescope removed from dynamic schedule by Jamie and Lucia (Ross)
- 11:18UT Telescope added back to dynamic schedule, bunch of earlier scans recorded before current time reached in schedule (Ross)
- 20:31UT Telescope removed for DS for APOD tracking.
- 21:36UT Telescope back in DS
Katherine
Disk VSN: HOB+0094
Volume at beginning: 0.000 GB
- 01:23UT - Tracking started after scheduler issues (Ross)
- 09:20UT Telescope removed from dynamic schedule by Jamie and Lucia (Bryn)
- 10:30UT Telescope added back to dynamic schedule (Bryn)
- 10:34UT +1day Telescope removed from dynamic schedule by Jamie and Lucia (Ross)
- 11:18UT Telescope added back to dynamic schedule (Ross)
- 15:19UT ALARM: error m5 -900 not while recording or playing. Once off, nothing seems wrong, no action was taken on previous occasions of this error occurring (previous dynamic observing experiments).
- 16:26UT ALARM: error m5 -900 not while recording or playing
- 20:31UT Telescope removed for DS for APOD tracking.
- 21:36UT Telescope back in DS
- 22:25UT Telescope removed for DS for APOD tracking.
- 23:00UT Telescope back in DS
Bryn Yarragadee
Disk VSN: HOB+0104
Volume at beginning: 0.000 GB
- 01:23UT - Tracking started after scheduler issues (Ross)
- 10:34UT +1day Telescope removed from dynamic schedule by Jamie and Lucia (Ross)
- 11:18UT Telescope added back to dynamic schedule, bunch of earlier scans recorded before current time reached in schedule (Ross)
- 22:25UT Telescope removed for DS for APOD tracking.
- 23:00UT Telescope back in DS
~~DISCUSSION~~